OpenStudy (pokemon23):

write each number as a power of 10 using negative exponents.

OpenStudy (pokemon23):

1/1000

OpenStudy (anonymous):

10^-3 i think

OpenStudy (pokemon23):

wouldn't this be 1/10-3

OpenStudy (pokemon23):

MIMI any suggestion?

OpenStudy (mimi_x3):

Yeah, its 10^-3

OpenStudy (anonymous):

it is 10^-3, cause that =1/10^3=1/1000